OpenStudy (anonymous):

please explain with step by step process! :) simplify ((x+1)/(x(x-1)))-((1)/(2(x-1)))

OpenStudy (astrophysics):

\[\frac{ (x+1) }{ x(x-1) }-\frac{ 1 }{ 2(x-1) }\] is this your problem?

OpenStudy (astrophysics):

Use the following rule and simplify. \[\large \frac{ a }{ b }-\frac{ c }{ d } \implies \frac{ ad-bc }{ bd }\]

OpenStudy (akashdeepdeb):

\[\frac{(x+1)}{x(x-1)} - \frac{1}{2(x-1)}\] Start by taking something common, out. \[\frac{1}{(x-1)}*(\frac{(x+1)}{x} - \frac{1}{2})\] Now, do what @Astrophysics said and simplify.